OpenStudy (anonymous):

Write the equation in logarithmic form. 1254/3 =625

OpenStudy (aravindg):

hey the question is wrong!

OpenStudy (aravindg):

1254/3=418

OpenStudy (anonymous):

@AravindG thats what the question asked was to put that into log form.

OpenStudy (aravindg):

but whatever form we put this is a wrong equation i nevr can agree 1=2 even if u take log on both sides

OpenStudy (anonymous):

yup AravindG is correct

OpenStudy (anonymous):

oh t h correct question is (125)^4/3=625 log representation i s log125 (625) =4/3 here base of log is 125

OpenStudy (anonymous):

@singerx13 u forgot that 4/3 is the index of 125

OpenStudy (dumbcow):

in general the rule is: \[\large b^{x} = y \leftarrow \rightarrow x = \log_{b} y \]